Hot Water Heater Due in 7 hours, 27 minutes A family comes home from a long vacation with laundry to do and showers to take. The water heater has been turned off during vacation. If the water heater has a capacity of 51.8 gallons and a 4610 W heating element, how much time is required to raise the temperature of the water from 22.0°C to 64.9°C? Assume that the heater is well insulated and no water is withdrawn from the tank during this time. Submit Answer Tries 0/10
Answer #1

1 gallon=3.785 litre

51.8*3.785=196.063 litre=196.063 kg

Q=m*c*delT

4610*t=196.063*4184*(64.9-22)

t=7633.85 seconds

=2hours 7minutes 12seconds